# PALM: PAnoramic Learning Map Integrating Learning Analytics

### v1

@@ II-B User Interface
[…]
@@ III-B System Evaluation (RQ2)
Figure 5: Screenshots of the existing systems (traditional curriculum map |||| attendance confirmation in LMS |||| web syllabus)
@@
[…]
@@ IV-A Effectiveness Evaluation (RQ1)
Intention 5.1 (1.07) 5.7 (0.83) $-4.1^{***}$ 0.77
Attitude 5.0 (0.67) 5.6 (0.90) $-4.3^{***}$ 0.80
Subjective norm 3.9 (1.32) 4.4 (1.47) $-3.1^{~{}**}$ 0.58
Behavioral control 4.1 (1.10) 5.3 (0.87) $-6.5^{***}$ 1.21
@@ IV-B System Evaluation (RQ2)
Visual attraction 3.4 (1.05) 6.0 (0.60) $-12.9^{***}$ 2.40
Usability 3.7 (1.07) 5.9 (0.77) ${}-9.9^{***}$ 1.84
Understanding level 3.8 (1.07) 6.1 (0.74) ${}-9.3^{***}$ 1.72
Perceived usefulness 4.0 (1.15) 5.7 (0.80) ${}-7.5^{***}$ 1.39
Behavioral changes 3.7 (1.15) 5.3 (0.87) ${}-6.9^{***}$ 1.29
@@
[…]

### v2

@@ II-B User Interface
  3. 02
    Individual Learning Engagement (Layer 2): The left half of each course block is overlaid with a blue-shaded area representing individual learning engagement, which is defined as the average of attendance rates, quiz scores, and assignment submissions. The combination of displayed information can be freely customized through display settings modal (A in Figure 3).
  4. 03
@@ III-B System Evaluation (RQ2)
Figure 5: Screenshots of the existing systems (traditional curriculum map $|$ attendance confirmation in LMS $|$ web syllabus)
@@
In addition to statistical analysis, qualitative data were collected through supplementary survey questions. Participants were asked to evaluate whether 01 course relevance lines in PALM aligned with their intuition, using a five-point Likert scale. To obtain more in-depth insights, open-ended questions solicited feedback on various UI components, including the curriculum map, course blocks, and engagement visualizations. Participants also provided suggestions for additional features or information they would like to see incorporated into PALM, as well as potential use cases they envisioned for the system.
@@ IV-A Effectiveness Evaluation (RQ1)
Intention 5.1 (1.07) 5.7 (0.83) $4.1^{***}$ 0.77
Attitude 5.0 (0.67) 5.6 (0.90) $4.3^{***}$ 0.80
Subjective norm 3.9 (1.32) 4.4 (1.47) $3.1^{~**}$ 0.58
Behavioral control 4.1 (1.10) 5.3 (0.87) $6.5^{***}$ 1.21
@@ IV-B System Evaluation (RQ2)
Visual attraction 3.4 (1.05) 6.0 (0.60) $12.9^{***}$ 2.40
Usability 3.7 (1.07) 5.9 (0.77) $9.9^{***}$ 1.84
Understanding level 3.8 (1.07) 6.1 (0.74) $9.3^{***}$ 1.72
Perceived usefulness 4.0 (1.15) 5.7 (0.80) $7.5^{***}$ 1.39
Behavioral changes 3.7 (1.15) 5.3 (0.87) $6.9^{***}$ 1.29
@@
The results of the qualitative evaluation were as follows. Regarding the 01 course relevance lines, excluding two participants who responded “I don’t know,” 89% of the remaining respondents found them “very intuitive” or “somewhat intuitive.” Supporting comments included: “the lines connect the recommended prerequisite courses” and “they reflect relationships I perceived when taking the courses.” However, some noted the lines were “not entirely convincing,” citing examples such as the connection between Compiler and Mathematics for Electrical Engineering as counterintuitive. These results suggest the TF-IDF based visualization was perceived as reasonably valid, though not universally intuitive.
Open-ended responses also provided broader impressions of the system. Compared to the traditional spreadsheet-style curriculum map, which was criticized for “small font size and poor readability,” the digitalized version was praised for its “improved text readability,” “clearer course relationships,” and “intuitive zoom functionality.” Additionally, features such as 02 visualizations of learning engagement and 03 comparisons with past course participants were positively received. Students noted that “the intensity of colors in course blocks facilitated intuitive comparisons” and that “academic grade could be linked to specific behaviors like attendance.” However, areas for improvement were also identified, including “excessive use of blue making elements hard to distinguish,” “unclear indicator meanings,” and the need for “the distribution and percentiles of grades.”